MICROWAVE SPECTROSCOPY OF PLATINUM MONOCYANIDE, PtCN

\underline{EMI Y. OKABAYASHI}, TOSHIAKI OKABAYASHI and MITSUTOSHI TANIMOTO, Department of Chemistry, Faculty of Science, Shizuoka University, Oya 836, Shizuoka.

The rotational spectrum of platinum monocyanide, PtCN, was observed by employing a source-modulation microwave spectrometer. The PtCN species was generated in a dc glow discharge through the mixture of CH_{3}CN and Ar by a sputtering reaction with a platinum sheet on a cathode. Paramagnetic lines were observed every 6GHz and assigned to three isotopomers, ^{194}PtCN, ^{195}PtCN, and ^{196}PtCN. There was no Λ-type doubling, but hyperfine splitting due to 195Pt nucleous for ^{195}PtCN. The hyperfine structure could be fitted to either ^{2}Π_{3/2} or ^{2}Δ_{5/2} case(c) Hamiltonian within experimental error. The nuclear - spin interaction constant C_{I} was derived to be around 0.2 MHz, which was one order of magnitude larger than that of ^{195}PtCO (C_{I} = 0.0242 MHz)^{a}. This result implies that low-lying electronic states would exist comparatively near to the ground electronic state, as in the case of NiCN^{b}^{c}.

aC. J. Evans and M. C. L. Gerry J. Phys. Chem. A \underline{105}, 9659 (2001).

bP. M. Sheridan and L. M. Ziurys J. Chem. Phys. \underline{118}, 6370 (2003).

cC. T. Kingston, A. J. Merer and T. D. Varberg J. Mol. Spectrosco. \underline{215(1)}, 106 (2002).